### v2

@@ 4 Experiments
Experiment Setup
For comprehensive experiments, we evaluate both close-sourced and open-sourced models. The closed-source models include the GPT family (o3-mini, o1-mini, GPT-4.1-mini, GPT-4o-mini and GPT-4o) [19, 17], Claude-3.5-Sonnet and Gemini-2.0-flash. The open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct, Qwen-2.5-72B-Instruct and QwQ-32B) [16, 37, 34], the DeepSeek family (DeepSeek-v3 and DeepSeek-R1) [27, 11], the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct) [8] and Yi-Coder-9B-Chat [30]. To eliminate potential data leakage and unnecessary evaluation overhead, we select the most recent 1000 problems from CodeFlowBench for tests. Each model is evaluated in both multi-turn and single-turn scenarios, with Pass@k and APD@k as the primary metrics. Please refer to Appendix D.1 and D.2 for more implementation details.
4.1 Main Experiments
Table 1: Performance comparison on CodeFlowBench with Pass@1 and APD@1 (Average Pass Depth) metrics. As a reference for overall APD, CodeFlowBench’s average problem depth is 1.94.
Model Pass@1 APD@1 (Average Pass Depth)
@@ 4.1 Main Experiments
Gemini-2.0-flash 17.3 31.6 0.496 0.183 0.527 0.782 1.042
Claude-3.5-Sonnet 13.6 36.0 0.414 0.117 0.444 0.648 1.167
Open-Source (7B-Level)
@@
Yi-Coder-9B-Chat 5.0 13.8 0.228 0.021 0.270 0.388 0.500
Open-Source (32B-Level)
@@
QwQ-32B-Preview 17.3 32.7 0.487 0.261 0.539 0.570 1.042
Open-Source (70B-Level)
@@
Open-Source (Large Model)
@@
Overall Performance
Table 1 presents the extensive experiments across 1,000 latest problems. The highest observed Pass@1 score of 24.4% and maximum APD@1 of 0.602 demonstrate the benchmark’s rigorousness, with even state-of-the-art LLMs struggling to achieve satisfactory performance. While closed-source models generally outperform the open-source ones, exceptions exist—notably DeepSeek-R1 surpasses several proprietary models. Notably, two model categories exhibit parameter-efficient competence: specialized coding models and reasoning models. Small coding models like Qwen2.5-Coder and Yi-Coder deliver surprising Pass@1 relative to model size, while reasoning models like QwQ-32B-Preview approach GPT-4o’s performance via test-time scaling. These findings suggest that domain-specific and reasoning models are promising directions for future advancements.
@@
Table 1 also reveals the substantial performance gap between the multi-turn and single-turn scenarios. This gap underscores the inherent complexity of multi-turn code generation, which requires models to "look before and after" for long-context coherence across iterative function implementations. Furthermore, we observe distinct features of models’ performance across single-turn and multi-turn scenarios: while models like Claude-3.5-Sonnet and DeepSeek-R1 excel in single-turn scenarios, they exhibit up to +60% performance degradation in multi-turn settings. In contrast, models such as GPT-4o and Llama-3.3-70B-Instruct demonstrate more consistent performance. The fundamental difference lies in the cognitive demands: single-turn evaluation examines complex reasoning capabilities, while multi-turn evaluation needs to generate reusable, modular functions, and integrate them into larger components. These insights expose under-explored limitations such as iterative development and dependency management.
Table 2: Performance comparison on CodeFlowBench with the APT@1 (Average Pass Turn) metric. As a reference for overall APT, CodeFlowBench’s average problem turn is 2.20.
Model Overall Turn1 Turn2 Turn3 Turn4 Turn5
@@
o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
Open-Source (7B-Level)
Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
Open-Source (32B-Level)
Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
Open-Source (70B-Level)
Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
Open-Source (Large Model)
Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
@@
The performance stratification across dependency depths reveals distinct model characteristics. Small open-source models (7B-32B parameters) show pronounced depth-specific biases—Qwen2.5-Coder-7B achieves 0.750 APD@1 on Depth4 problems versus 0.018 on Depth1, suggesting emergent compositional abilities despite limited parameter counts. In contrast, larger models exhibit more balanced performance profiles, with GPT-4.1-mini maintaining 0.265-1.042 APD@1 across depths. Surprisingly, top performers like DeepSeek-R1 show no significant advantage on Depth4 problems (APD@1: 0.916 vs. 1.208 for DeepSeek-V3), indicating current reasoning patterns inadequately handle deep dependency chains. This suggests models could benefit from explicit dependency tracking mechanisms and iterative verification loops during code generation. For another perspective, we report APT@1 in Table 2. Notably, the trends of APT closely align with APD, reinforcing our observations of models’ depth-wise performance.
@@ 4.2 Analysis and Discussion
A deeper look of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential function compositions). However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), state-of-the-art models exhibit significant performance degradation. This phenomenon is empirically validated in Figure 5, which illustrates the Pass@1 scores across varying turn counts. The consistent performance trajectory—initially high for 1-2 turn problems followed by approximately exponential decline as turn counts increase—demonstrates the inherent challenges in multi-turn code generation. Even top-performing models (e.g., Deepseek-R1 with 20.5% Pass@1) fail to solve problems requiring more than six turns. This underscores the critical limitation to balance local correctness and global integration across iterative development cycles.
Figure 5: The Pass@1 metrics on multi-turn problems grouped by model categories and turn number.
To quantity dependency structure complexity, we introduce the Dependency Structure Complexity (DSC) metric, defined as the ratio of total turns to the maximum depth in the AST. Figure 13 presents the models’ performance across different DSC intervals, revealing that most models perform well on problems with linear dependency structures but struggle significantly as the dependency structure becomes more complex. We leave the discussion on DSC in Appendix E.1.
@@
Given the significant performance gap between models in multi-turn and single-turn scenarios, we conducted studies to identify the underlying reasons. We categorized errors into three primary types: (1) Incomplete Reasoning (IR): Models often handle only straightforward "happy-path" cases and fail to generalize. They may oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapses on larger inputs. This reflects a limitation in the models’ reasoning abilities. (2) Insufficient Globalization (IG): While a function’s logic may run correctly in isolation, it may omit necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This indicates a limitation in the models’ ability to manage global context. (3) Instruction Misinterpretation (IM): Given multi-turn prompts, models could solve isolated subproblems but lack a coherent understanding of the overarching goal. Typical failures include misusing helper functions or implementing disorganized code within the top-level function, i.e. incorrect integration of components. To quantify the distribution of errors, we randomly sampled examples and manually annotated error categories. The proportion is calculated in Table 3.
Table 3: Proportion of each error type on Deepseek-V3 and o1-mini. The IR Remediation Rate is the proportion of IR errors in multi-turn scenario which are correctly answered in single-turn scenario.
Model IR(%) IG(%) IM(%) Others(%) IR Remediation Rate(%)
Deepseek-v3 47.2 14.6 32.6 5.6 16.7
o1-mini 38.7 11.8 43.0 6.5 8.3
Table 3 offers several insights. First, Incomplete Reasoning remains the dominant error type which is the main challenge of competition-level problems. To distinguish the effect of multi-turn scenario, we further introduce IR Remediation Rate to quantify how much of IR can be fixed when switching from multi-turn to single-turn scenarios. The statistics show that, most IR errors (30% of all error types) originate from the multi-turn scenario. Furthermore, Insufficient Globalization and Instruction Misinterpretation are inherently specific to multi-turn scenarios. The high rate of IM errors highlights that models struggle with consistent instruction following across multiple turns, or the codeflow task. Finally, IG errors reveal that models often lack "global awareness", highlighting the importance of both local correctness and global coherence evaluation. We leave case studies in Appendix E.2.
@@ 5 Conclusion
This paper introduces CodeFlowBench, the first benchmark specifically designed to evaluate multi-turn iterative code generation capabilities in realistic development workflows, i.e., the codeflow. Comprising 5,258 competition-level programming problems curated from Codeforces, our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APT@k, APD@k, and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 60% performance degradation). Our fine-grained analysis identifies dominant failure modes and provides insights for further advancements. Extensive experiments across 16 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also paves the way for more realistic and powerful code generation systems. Future Work: CodeFlowBench currently remains confined to competition-level problems, which should be expanded with repository-level challenges. We also plan to design and train dedicated code generators tailored to the codeflow task.
@@
[…]
Appendix A Features of CodeFlowBench
We collect problems from Codeforces, which offer following advantages to our benchmark:
High difficulty.
Even state-of-the-art LLMs have low pass rates on Codeforces problems [32]. Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful.
High quality.
Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.
Periodic updates.
To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly. Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.
Appendix B Data Curation Pipeline
B.1 Detail of Problem Scraping

### v3

@@ 4 Experiments
Model Pass@1 (%) APD@1 (Average Pass Depth)
@@
Gemini-2.5-flash-thinking 22.7 39.8 0.642 0.167 0.761 0.981 1.208
Gemini-3-flash-thinking 48.4 65.5 1.120 0.445 1.303 1.625 1.625
Claude-4-Sonnet 27.3 42.9 0.830 0.185 0.893 1.250 1.375
Claude-4.5-Sonnet 31.6 47.9 0.857 0.222 1.008 1.377 1.521
Open-Source (7B)
@@
Qwen3-8B 9.1 18.3 0.406 0.088 0.444 0.738 1.125
Open-Source (32B)
@@
Qwen3-Coder-30B-A3B 11.6 26.0 0.405 0.100 0.458 0.681 0.917
Open-Source (70B)
@@
Open-Source (Large)
@@
CodeFlowBench-Repo
Closed-Source
GPT-5 34.7 48.8 0.898 0.483 1.020 1.385 3.200
GPT-5-mini 28.1 53.5 0.740 0.362 0.725 1.615 3.000
o3-mini 28.3 51.2 0.850 0.466 0.902 1.615 2.800
o1-mini 21.4 48.8 0.817 0.362 0.860 1.615 3.600
GPT-4.1-mini 27.6 48.0 0.724 0.431 0.705 1.692 1.800
GPT-4o-mini 24.6 43.3 0.667 0.414 0.800 1.000 1.400
Gemini-2.5-flash-thinking 26.1 44.6 0.765 0.375 0.833 1.167 3.600
Gemini-3-flash-thinking 33.9 56.7 0.905 0.500 1.020 1.308 3.400
Claude-4-Sonnet 33.9 57.6 0.850 0.461 0.980 1.385 2.600
Claude-4.5-Sonnet 33.1 56.7 0.913 0.448 1.039 1.615 3.200
Open-Source (7B)
Llama-3.1-8B-Instruct 17.3 39.6 0.417 0.310 0.451 0.615 0.800
Qwen2.5-Coder-7B-Instruct 19.7 39.1 0.528 0.310 0.647 0.692 1.400
Qwen3-8B 18.1 36.2 0.394 0.293 0.450 0.462 0.800
Open-Source (32B)
Qwen2.5-Coder-32B-Instruct 25.2 45.7 0.646 0.379 0.784 1.000 1.400
Qwen3-Coder-30B-A3B 23.6 42.5 0.543 0.414 0.470 1.385 0.600
Open-Source (70B)
Llama-3.3-70B-Instruct 22.1 39.6 0.614 0.345 0.522 1.846 1.400
Qwen2.5-72B-Instruct 25.2 43.3 0.716 0.397 0.745 1.230 2.800
Open-Source (Large)
Deepseek-V3 29.1 45.1 0.771 0.483 0.745 1.538 2.400
Deepseek-R1 29.4 52.0 0.770 0.431 0.760 1.538 2.800
Table 1: Performance on CodeFlowBench-Comp and CodeFlowBench-Repo across different model scales. Both datasets report Pass@1 and APD@1 (Average Pass Depth) across varying depths.
Experiment Setup
For comprehensive experiments, we evaluate both close-sourced and open-sourced models. The closed‑source models include the GPT (gpt5), Gemini (gemini3) and Claude families (claude4). The open‑source models include the Qwen (yang2025qwen3), DeepSeek (liu2024deepseek; guo2025deepseek) and Llama families (grattafiori2024llama). We experiments on CodeFlowBench-Comp and CodeFlowBench-Repo under identical settings. For CodeFlowBench-Comp, we selected a test subset comprising the most recent 1,000 problems to eliminate potential data leakage and decrease evaluation overhead while maintaining statistical significance. All models are evaluated in both multi-turn and single-turn scenarios, utilizing Pass@k and APD@k as the primary metrics. Please refer to Appendix E.1 and E.2 for implementation details.
4.1 Main Experiments
Table 1 presents the main experiment results. For the majority of LLMs, the Pass@1 in the multi-turn scenario remains below 30%, underscoring CodeFlowBench’s difficulty. The performance across CodeFlowBench also reveals divergent model specializations. For example, Gemini-3-flash performs best on CodeFlowBench-Comp while GPT-5 outperforms on CodeFlowBench-Repo. Furthermore, we present more analysis as follows.
@@ 4.1 Main Experiments
Table 1 reveals a substantial performance gap between multi-turn and single-turn scenarios across both CodeFlowBench-Comp and CodeFlowBench-Repo. This consistent degradation underscores the inherent complexity of multi-turn generation, which necessitates long-context coherence and robust dependency management. Notably, we observe a divergence in model behaviors: while reasoning-enhanced models like Deepseek-R1 and o3-mini achieve impressive single-turn scores, they suffer performance drops exceeding 50% in multi-turn settings. In contrast, models such as GPT-5 demonstrate greater stability with narrower gaps. This distinction suggests that single-turn metrics largely reflect peak reasoning capabilities, whereas multi-turn evaluation exposes previously under-explored limitations in iterative development and dependency management.
Model CodeFlowBench-Comp CodeFlowBench-Repo
Overall Turn 1 Turn 2 Turn 3 Turn 4 Overall Turn 1 Turn 2 Turn 3+
@@
GPT-5 0.832 0.185 0.951 1.224 1.781 0.890 0.509 1.049 1.000
GPT-5-mini 0.717 0.103 0.832 1.158 1.345 0.520 0.363 0.537 0.667
o3-mini 0.600 0.322 0.632 0.777 0.983 0.780 0.455 0.854 1.056
o1-mini 0.581 0.233 0.645 0.798 0.879 0.675 0.345 0.725 0.944
GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.701 0.436 0.732 1.056
GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.571 0.418 0.650 0.778
Gemini-2.5-flash-thinking 0.695 0.167 0.855 0.923 1.218 0.687 0.364 0.879 0.600
Gemini-3-flash-thinking 1.178 0.445 1.386 1.617 1.800 0.905 0.491 1.000 1.000
Claude-4-Sonnet 0.830 0.185 0.990 1.153 1.564 0.866 0.455 1.024 1.389
Claude-4.5-Sonnet 0.927 0.222 1.104 1.291 1.704 0.961 0.418 1.146 1.167
Open-Source (7B)
Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.370 0.327 0.439 0.389
Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.528 0.327 0.512 0.778
Qwen3-8B 0.452 0.075 0.490 0.692 1.036 0.362 0.309 0.488 0.333
Open-Source (32B)
Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.583 0.364 0.780 0.833
Qwen3-Coder-30B-A3B 0.449 0.100 0.508 0.699 0.836 0.496 0.418 0.561 0.333
Open-Source (70B)
Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.516 0.345 0.553 0.750
Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.669 0.364 0.780 0.889
Open-Source (Large)
Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.709 0.455 0.829 0.722
Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.675 0.418 0.925 0.611
Table 2: Pass Turn comparison across models on CodeFlowBench-Comp and CodeFlowBench-Repo. Consistent with Table 1, the overall performance distribution shows that models with high APD@1 scores also exhibit high APT@1 scores, validating the precision of our metric.
@@
Our analysis across dependency depths reveals a dichotomy in scaling laws. First, in shallow dependencies, reasoning paradigms supersede parameter scale. Efficient reasoning models, such as o3-mini, significantly outperform larger base models like GPT-4o on CodeFlowBench-Comp, suggesting that logical deduction is more critical than raw capacity for immediate or localized tasks. Second, deep dependencies function as a capacity filter, where model scale becomes decisive. In CodeFlowBench-Repo, Qwen2.5-72B-Instruct doubles the performance of its 7B and 32B counterparts, indicating that maintaining coherence across extended dependency chains requires substantial parameter capacity. Finally, SOTA models like Gemini-3-flash-thinking and GPT-5 demonstrate full-spectrum dominance, eliminating trade-offs by combining agile reasoning for local logic with the extensive capacity required for global integration.
From another perspective, we report APT@1 in Table 2. Notably, the APT trends closely align with APD, reinforcing our observations of models’ depth-wise performance.
@@ 4.2 Analysis and Discussion
A deeper analysis of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential compositions). However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), models exhibit significant performance degradation. This phenomenon is empirically validated in Figure 4, which illustrates the Pass@1 scores across varying turn counts. The consistent performance trajectory demonstrates the inherent challenges in multi-turn code generation, where curves initially high for 1-2 turn problems and followed by sharp decline as turn counts increase. Even top-performing models fail to solve problems requiring more than six turns. This underscores the critical limitation to balance local correctness and global integration across iterative development cycles.
Figure 4: Models’ Pass@1 results on multi-turn problems grouped by model categories and turn number.
To quantity dependency structure complexity, we introduce the Dependency Structure Complexity (DSC) metric, defined as the ratio of total turns to the maximum depth in the AST. Figure 16 presents the models’ performance across different DSC intervals, revealing that most models perform well on problems with linear dependency structures but struggle significantly as the dependency structure becomes more complex. For space reason, we leave more discussion in Appendix F.1.
@@
Given the significant performance gap between models in multi-turn and single-turn scenarios, we conducted studies to identify the underlying reasons. We categorized errors into three primary types: (1) Incomplete Reasoning (IR): Models often handle only straightforward "happy-path" cases and fail to generalize. They may oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapses on larger inputs. This reflects a limitation in the models’ reasoning abilities. (2) Insufficient Globalization (IG): While a function’s logic may run correctly in isolation, it may omit necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This indicates a limitation in the models’ ability to manage global context. (3) Instruction Misinterpretation (IM): Given multi-turn prompts, models could solve isolated subproblems but lack a coherent understanding of the overarching goal. Typical failures include misusing helper functions or implementing disorganized code within the top-level function, i.e. incorrect integration of components. To quantify the distribution of errors, we randomly sampled examples from CodeFlowBench-Comp and CodeFlowBench-Repo. The error categories are manually annotated by authors. The proportion is calculated in Table 3.
Model Dataset IR(%) IG(%) IM(%) Others(%)
DeepSeek-V3 Comp 47.2 14.6 32.6 5.6
Repo 32.1 22.6 44.0 1.2
o1-mini Comp 38.7 11.8 43.0 6.5
Repo 23.7 18.3 55.9 2.2
Gemini-3-Flash Comp 52.3 20.5 27.3 0.0
Repo 34.6 21.8 41.0 2.6
Table 3: Error distribution across models and benchmarks. Values indicate the percentage of each error category relative to the total number of errors identified in the sampled dataset for each model.
Table 3 offers several insights. First, the distinct error distributions confirm that our benchmark components target divergent primary challenges: CodeFlowBench-Comp is dominated by IR and CodeFlowBench-Repo emphasizes global context which shifting the burden to IG and IM. Second, distinct bottlenecks constrain models of different tiers. DeepSeek-V3 and o1-mini are primarily limited by IM, suggesting their performance is capped at instruction comprehension and adherence. In contrast, for SOTA models like Gemini-3-Flash, the bottleneck shifts back to IR, indicating task understanding becomes minor issues. Detailed case studies are provided in Appendix F.2.
@@ 5 Conclusion
This paper introduces CodeFlowBench, the first benchmark specifically designed to evaluate multi-turn iterative code generation capabilities in realistic development workflows, i.e., the codeflow. CodeFlowBench consists of two subsets: CodeFlowBench-Comp for multi-step algorithmic reasoning via competition challenges, and CodeFlowBench-Repo for evaluating generalization in real-world software contexts. Our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APD@k and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 50% performance degradation). Our fine-grained analysis identifies dominant failure modes and provides insights for further advancements. Extensive experiments across 19 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also paves the way for more realistic and powerful code generation systems.
Limitations
We acknowledge two limitations that present opportunities for future expansion. First, while the pipeline’s subproblem generation phase is general, the data preparation phase is currently specific. To achieve universal compatibility, we can consider implementing an agentic framework for automatically adapting diverse codebases. This will expand our pipeline’s coverage to a broader range of data sources beyond Codeforces and GitHub, such as Bitbucket, establishing a truly source-agnostic, end-to-end framework. Second, we currently lack models specifically fine-tuned for the codeflow task. Future work will consider fine-tuning specialized models and agentic systems on this data to assist developers in real-world, iterative programming scenarios. We also aim to extend our codeflow paradigm to include iterative debugging loops. By training specialized systems on these dynamic scenarios, we seek to enable robust and sustainable code implementation and maintenance, allowing models to effectively resolve requirements through multi-turn refinement.
Ethical Consideration
Codeforces provides an official “Codeforces materials usage license” [^2] https://codeforces.com/blog/entry/967. We have taken great care to make our pipeline fully comply with all licensing requirements and community guidelines:
Firstly, all problems we use are officially published by Codeforces, and we provide clear attribution to Codeforces. Our usage is strictly non-commercial and academic, in accordance with the license’s permitted use cases. Our data pipeline strictly avoids scraping any user-submitted code, forum discussion content or third-party solutions. We only collect officially published materials, namely problem statements and editorials.
Secondly, officially published Codeforces materials have long been used for widely recognized code generation benchmarks such as CodeContest (li2022competition), LiveCodeBench (jain2024livecodebench) and CodeElo (quan2025codeelo). Our work follows this community tradition and with the utmost respect for Codeforces.
Besides, we will also make our code properly hosted and clearly state its intended use to ensure responsible data collection.
@@
Appendix A Data Curation Pipeline
A.1 Detail of Problem Scraping